# Socially Aware Robot Navigation through Scoring Using Vision-Language Models

### v1

@@ III-B VLM-based Social Navigation Architecture
Figure 3: An example input image ($\mathcal{I}^{t}$) and prompt ($\mathcal{P}$) used in VLM-Social-Nav. Parameterized inputs ($\mathbf{a}^{t}$) are highlighted in blue. Formatted outputs are highlighted in red. The example input data is one of the frontal approach scenarios from MuSoHu [nguyenmusohu]. The output of VLM is Move right with slow down, to slowly pass by a person on the right side, which was the same action the collected data took.
@@ III-C VLM-based Scoring Module
where $\mathcal{I}^{t}$ is an RGB image of the robot view at time $t$, $\mathcal{P}$ is a textual prompt, $\mathbf{a}^{t}$ is a current robot action at time $t$. Inspired by In-Context Learning (ICL), our prompt $\mathcal{P}$ is designed to leverage the VLM’s reasoning abilities through zero-shot examples. This approach offers an interpretable interface, mirroring human reasoning and decision-making processes, without extensive training [min2022rethinking].
Our VLM-based scoring module starts from the insight that the action space of a mobile robot can be readily mapped to linguistic terms. For example, the action “move forward at a constant speed” can be linked to a linear velocity of $v^{t}$ m/s and an angular velocity of 0. The heading direction on the left indicates a positive value of $w^{t}$, while the direction on the right indicates a negative. Leveraging this understanding, we structure the output of the VLM into a linguistic format comprising the heading and the speed. Subsequently, our scoring module extracts $\mathcal{B}_{h}^{t+1}\mapsto(v_{h}^{t+1},w_{h}^{t+1})\in\mathcal{A}$ from these tokens; $v_{h}^{t+1}=v^{t}+\delta_{s}$, where $\delta_{s}$ is derived from the response for SPEED; $w_{h}^{t+1}=\delta_{d}$, where $\delta_{d}$ is derived from the response for DIRECTION. Thus, the social cost term for the next time step can be calculated:
$C_{social}^{t+1}=w_{l}\cdot\|v-v_{h}^{t+1}\|+w_{a}\cdot\|w-w_{h}^{t+1}\|,$ (4)
where $w_{l}$ and $w_{a}$ are a non-negative weight values. Given all the cost terms, our low-level optimization-based motion planner finds the robot action $(v^{*},w^{*})$ that minimizes the cost.
@@
Figure 4: Qualitative Results: the robot motion results with VLM-Social-Nav for four social navigation scenarios: (a) Frontal Approach, (b) Frontal Approach with Gesture, (c) Intersection, and (d) Narrow Doorway. The gray solid arrow shows the participant’s path. The red solid arrow shows the robot’s path. The red and gray dashed arrows show the robot’s and participant’s path respectively after a stop motion. A caption on the top left shows the result from the VLM.
TABLE I: Quantitative Results: performance comparisons using BC [pomerleau1988alvinn] method trained on SCAND dataset [karnan2022scand], DWA [fox1997dynamic], and VLM-Social-Nav. We achieve at least $36.37\%$ improvement in average success rate and $20.00\%$ improvement in average collision rate in the four social navigation scenarios. The user study score shows that VLM-Social-Nav generates the most socially compliant navigation behavior.
@@
Collision Rate (%) BC 46.66 60.00 33.33 40.00
DWA 26.66 20.00 13.33 33.33
VLM-Social-Nav 6.66 0 0 6.66
@@ IV-A Implementation Details
VLM-Social-Nav is tested on a Turtlebot 2 equipped with a Velodyne VLP16 lidar, and a Zed 2i camera, and a laptop with intel i7 CPU and an Nvidia GeForce RTX 2080 GPU. We use YOLO [YOLO2023] as our real-world perception model to detect key objects (e.g., humans, door, and gesture) and Generative Pre-trained Transformer 4 with Vision (GPT-4V) [gpt4v2023] as our VLM to comprehend the social dynamics and output the immediate preferred robot action. We combined our method with a low-level motion planner DWA [fox1997dynamic]. We compare VLM-Social-Nav with DWA without social cost $C_{social}$ and BC [pomerleau1988alvinn] method trained on a state-of-the-art, large-scale social navigation dataset, SCAND [karnan2022scand].
To validate VLM-Social-Nav, we carefully follow the social robot navigation studies [francis2023principles, pirk2022protocol] that have set up the benchmark scenarios and the metrics for measuring social navigation. We present qualitative, quantitative, and user study results in four different indoor social navigation scenarios:
@@ IV-B Qualitative Result
  • Frontal Approach: The robot is expected to yield or slow down and modify its original trajectory so as not to obstruct the human path. Similar to driving rules in North America, it is conventional to keep on the right side.
@@ IV-C Quantitative Result
Table I reports the results averaged over 15 runs for each method and scenario. The results demonstrate that VLM-Social-Nav, DWA with social cost, outperforms other methods in every metric. DWA excels at following a path smoothly, yet it faces challenges in collision avoidance when the participant walks too quickly and the robot attempts to make an abrupt change in motion. We also observed that the outcomes of BC varied. At times, when attempting to avoid collisions, it failed to return to its original path and failed to reach the goal. Conversely, there were instances where it didn’t attempt collision avoidance at all, resulting in collisions with the participants. For gesture recognition, only our proposed method successfully responded to the participants’ gestures. In total, we achieve at least $36.37\%$ improvement in the average success rate and $20.00\%$ improvement in the average collision rate in the four social navigation scenarios.
@@ IV-D User Study
To validate the social compliance of VLM-Social-Nav, we conduct a user study. We ask the participants to walk along the predefined trajectory and ask them to answer questionnaires about the robot motion [pirk2022protocol] (Table II). * denotes negatively formulated questions, for which we reverse-code the ratings to ensure comparability with the positively formulated ones. The three methods are randomly shuffled and repeated three times. Each scenario is tested on five different participants. We use a five-level Likert scale to ask participants to rate their agreement toward these statements.
Fig. 5 and the user study score on Table I show the study result. The plot shows the per-question average scores for the three methods in each scenario. Based on the results, it’s evident that VLM-Social-Nav received the highest level of agreement from participants across all questions, indicating its strong adherence to social norms. The standard error of the BC method was also large, indicating that the performance of the BC method was not constant. The score difference between VLM-Social-Nav and DWA was not large in the narrow doorway scenario. This is because, when attempting to enter the narrow doorway, the robot motion results from DWA indicated that it failed to find a plan and froze, resembling almost a complete stop.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ V Conclusion
We propose a novel social navigation approach based on Vision-Language Models (VLMs), focusing on real-time, socially compliant decision-making in human-centric environments. We utilize the perception model to detect important social entities and prompt a VLM to generate guidance for socially compliant behavior. VLM-Social-Nav features a VLM-based scoring that ensures socially appropriate and effective robot actions. This minimizes the dependence on extensive training datasets and eliminates the necessity for explicit rules or rewards typically associated with reinforcement learning methods. By furnishing textual instructions to VLM, we can instruct the robot to adhere to specific navigation rules, such as navigating on the right or left according to cultural norms. In practice, it results in improved socially compliant navigation in human-shared environments. We demonstrate and evaluate our system in four different real-world social navigation scenarios with a Turtlebot robot.
Although our robot experiments were conducted only indoors, according to the example in Fig. 3, VLM-Social-Nav can be extended to outdoor scenarios. Our immediate focus will be on advancing our work into a global outdoor navigation system. VLM-Social-Nav is taking certain social entities to query VLM. To make it work on more general cases, when or where to query VLM will be a critical consideration. Our experiments revealed that VLM excels when presented with choices rather than being tasked with making singular, decisive actions. We aim to explore more on the methods to provide VLM with better navigation options beyond simple direction and speed adjustments.

### v2

@@ 17       Update robot position $\mathbf{p}_{r}$;
Figure 3: An example input image ($\mathcal{I}^{t}$) and prompt ($\mathcal{P}$) used in VLM-Social-Nav. Parameterized inputs ($\mathbf{a}^{t}$) are highlighted in blue. Formatted output specifying the heading ($\delta_{d}$) and the speed ($\delta_{s}$) are highlighted in red. The example input data is one of the frontal approach scenarios from MuSoHu [nguyenmusohu]. The output of VLM is Move right with slow down, which was the same action the collected data took.
@@ III-C VLM-based Scoring Module
where $\mathcal{I}^{t}$ is an RGB image from the robot view at time $t$, $\mathcal{P}$ is a textual prompt, $\mathbf{a}^{t}$ is a current robot action at time $t$. Inspired by In-Context Learning (ICL), our prompt $\mathcal{P}$ is designed to leverage the VLM’s reasoning abilities through zero-shot examples. This approach offers an interpretable interface, mirroring human reasoning and decision-making processes, without extensive training [min2022rethinking].
Our VLM-based scoring module starts from the insight that the action space of a mobile robot can be readily mapped to linguistic terms. For example, the action “move forward at a constant speed” can be linked to a linear velocity of $v^{t}$ m/s and an angular velocity of 0. The heading direction on the left indicates a positive value of $w^{t}$, while the direction on the right indicates a negative value. Leveraging this understanding, we structure the output of the VLM into a linguistic format comprising the heading and the speed. Subsequently, our scoring module extracts $\mathcal{B}_{h}^{t+1}\mapsto(v_{h}^{t+1},w_{h}^{t+1})\in\mathcal{A}$ from these tokens; $v_{h}^{t+1}=v^{t}+\delta_{s}$, where $\delta_{s}$ is derived from the response for the speed; $w_{h}^{t+1}=\delta_{d}$, where $\delta_{d}$ is derived from the response for the heading. Thus, the social cost term for the next time step can be calculated:
$\mathcal{C}_{\textrm{social}}^{t+1}=w_{l}\cdot\|v-v_{h}^{t+1}\|+w_{a}\cdot\|w-w_{h}^{t+1}\|,$ (4)
where $w_{l}$ and $w_{a}$ are non-negative weight values. Given all the cost terms, our low-level optimization-based motion planner finds the robot action $(v^{*},w^{*})$ that minimizes the cost.
@@
Figure 4: Qualitative Results: the robot navgation behaviors with VLM-Social-Nav for four social navigation scenarios: (a) Frontal Approach, (b) Frontal Approach with Gesture, (c) Intersection, and (d) Narrow Doorway. The gray solid arrow shows the participant’s path. The red solid arrow shows the robot’s path. The red and gray dashed arrows show the robot’s and participant’s path respectively after a stop motion. A caption on the top left shows the result from the VLM.
TABLE I: Quantitative Results: performance comparisons using BC [pomerleau1988alvinn] trained on SCAND dataset [karnan2022scand], DWA [fox1997dynamic], and VLM-Social-Nav. VLM-Social-Nav achieves at least $27.38\%$ improvement in average success rate and $19.05\%$ improvement in average collision rate in the four social navigation scenarios. The user study score shows that VLM-Social-Nav generates the most socially compliant navigation behavior.
@@
Collision Rate (%) BC 42.86 66.67 28.57 38.10
DWA 28.57 19.05 19.05 38.10
VLM-Social-Nav 14.29 0 4.76 9.52
@@ IV-A Implementation Details
VLM-Social-Nav is tested on a Turtlebot 2 equipped with a Velodyne VLP16 LiDAR, and a Zed 2i camera, and a laptop with intel i7 CPU and an Nvidia GeForce RTX 2080 GPU. We use YOLO [YOLO2023] as our real-world perception model to detect key objects (e.g., humans, door, and gesture) and Generative Pre-trained Transformer 4 with Vision (GPT-4V) [gpt4v2023] as our VLM to comprehend the social dynamics and output the immediate preferred robot action. We combined our method with a low-level motion planner DWA [fox1997dynamic]. We compare VLM-Social-Nav with DWA without social cost $\mathcal{C}_{\textrm{social}}$ and BC [pomerleau1988alvinn] method trained on a state-of-the-art, large-scale social navigation dataset, SCAND [karnan2022scand]. The dataset contains various examples of socially compliant navigation behaviors teleoperated by humans including sticking to the right of the road, waiting for a human to pass, and following a crowd.
To validate VLM-Social-Nav, we carefully follow the social robot navigation studies [francis2023principles, pirk2022protocol] that have set up the benchmark scenarios and the metrics for measuring social compliance. We present qualitative, quantitative, and user study results in four different indoor social navigation scenarios:
@@ IV-B Qualitative Result
  • Frontal Approach: The robot is expected to yield or slow down and modify its original trajectory not to obstruct the human path. Similar to driving rules in North America, it is conventional to keep on the right.
@@ IV-C Quantitative Result
Table I reports the results averaged over 21 runs for each method and scenario. The results demonstrate that VLM-Social-Nav, DWA with social cost, outperforms other methods in every metric. DWA excels at following a path smoothly, yet it faces challenges in collision avoidance as it relies solely on the LiDAR sensor and it does not consider socially compliance. Most of the collisions occurred when DWA navigated in a way that interfered with a person’s path, for example, going in front of the person when intersecting. We also observe that the outcomes of BC varied. At times, when attempting to avoid collisions, it failed to return to its original path and failed to reach the goal. Conversely, there were instances where it didn’t attempt collision avoidance at all, resulting in collisions with the participants. For gesture recognition, only our proposed method successfully responded to the participants’ gestures. In total, VLM-Social-Nav achieves at least $27.38\%$ improvement in the average success rate and $19.05\%$ improvement in the average collision rate in the four social navigation scenarios.
@@ IV-D User Study
To validate the social compliance of VLM-Social-Nav, we conduct a user study. We ask the participants to walk along the predefined trajectory and ask them to answer questionnaires about the robot motion [pirk2022protocol] (Table II). * denotes negatively formulated questions, for which we reverse-code the ratings to ensure comparability with the positively formulated ones. The three methods are randomly shuffled and repeated three times. Each scenario is tested on seven participants. We use a five-level Likert scale to ask participants to rate their agreement toward these statements.
Fig. 5 and the user study scores on Table I show the study result. The plot shows the per-question average scores for the three methods in each scenario. Based on the results, it’s evident that VLM-Social-Nav receives the highest level of agreement from participants across all questions, indicating its strong adherence to social norms. The standard error of the BC method was large, indicating that the performance of the BC method was not consistent. The score difference between VLM-Social-Nav and DWA was not large in the narrow doorway scenario. This is because, when attempting to enter the narrow doorway, the robot motion results from DWA indicated that it failed to find a plan and froze, resembling almost a complete stop as the result of VLM-Social-Nav.
@@
IV-E Discussion
Real-time navigation with VLM: GPT-4 and similar large VLMs require several seconds to respond to prompts, making continuous querying impractical for real-time navigation tasks. To address this, we optimized VLM-Social-Nav in two ways: first, by formatting prompts and providing predefined choices, which resulted in reduced response times. Second, we minimize queries by using a perception model to detect social cues, allowing for timely VLM queries only when necessary. These choices enable and allow average response times of 2-3 seconds, sufficient for human interaction and navigation. While such a limitation can be problematic in more dynamic scenarios that require frequent interactions, future advancements in fast large language models promise further extensions of our approach.
Socially aware navigation with VLM: We observe that VLMs can analyze and reason about social interactions from single images. Using various single images, including those collected by ourselves and from social robot navigation datasets [nguyenmusohu, karnan2022scand], VLMs accurately describe scenes and suggest socially compliant navigation strategies with reasons. For instance, with the image shown in Fig. 3, GPT-4V describes the scene as a person is walking towards the camera along a sidewalk. To navigate this situation, GPT-4V advises to yield the right of way because it is generally customary to keep to the right side of the path when encountering someone coming from the opposite direction, similar to driving rules. However, despite their powerful capabilities, VLMs can still make mistakes. Therefore, it is not yet safe to rely solely on VLMs for navigation decisions. Instead, we incorporate their output as a cost term in our overall decision-making process.
A single action output of VLM: Our approach begins with the observation that the robot action can be effectively translated into linguistic terms. VLMs still face challenges in accurate spatial understanding. Mapping spatial information from images to a textual format for communicating with VLMs is another challenging research problem, unless we use a fine-tuned foundation model. Therefore, adopting an approach where robot actions can be easily translated into textual output offers a straightforward yet robust solution that effectively mitigates VLM hallucination. However, this may have the drawback of not being able to plan long-range trajectories ahead. This is our immediate interest for future work; to extend our approach to planning longer-range trajectories for outdoor navigation.
More challenging scenarios: Although our robot experiments were conducted only indoors, according to the example in Fig. 3, VLM-Social-Nav can be extended to outdoor scenarios in more complex environments. VLMs successfully retrieve environment information that is significant for outdoor social robot navigation, such as sidewalks, zebra crossings, and cars. This will be our immediate focus for future work, i.e., to advance our work into global outdoor navigation. We also aim to extend our approach to complex scenarios involving multiple individuals. When tested with the scenario with a group of people, VLM-Social-Nav successfully outputs socially compatible actions, realizing the group of people in the scene. However, when multiple groups are present, simple directions like left or right may not suffice to describe effective robot navigation. We plan to explore multiple trajectory generation methods [liang2023mtg, liang2024dtg] and selection techniques using visual prompts [nasiriany2024pivot, liu2024moka] to enhance interaction precision.
@@ V Conclusion
We propose a novel social navigation approach based on VLMs, focusing on real-time, socially compliant decision-making in human-centric environments. We utilize the perception model to detect important social entities and prompt a VLM to generate guidance for socially compliant behavior. VLM-Social-Nav features a VLM-based scoring that ensures socially appropriate and effective robot actions. This minimizes the dependence on extensive training datasets and eliminates the necessity for explicit rules or rewards typically associated with reinforcement learning methods. By furnishing textual instructions to VLM, we can instruct the robot to adhere to specific navigation rules, such as navigating on the right or left according to cultural norms. In practice, it results in improved socially compliant navigation in human-shared environments. We demonstrate and evaluate our system in four different real-world social navigation scenarios with a Turtlebot robot.